OpenStudy (anonymous):

Hockey teams receive 2 points for a win and 1 point for a tie. The Wildcats once won a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the Wildcats have?

OpenStudy (nurali):

This is a system of equations problem: Let w stand for wins, and t stand for ties, then we have: (2*w)+(1*t)=60 given w=t+9 given Let's use substitution, using the value of w in Equation 1 in Equation 2: 2*(t+9)+t=60 2t+18+t=60 3t+18=60 3t=42 t=14 Use this value of t in Equation 2 to find w: w=t+9 w=14+9 w=23 So on their championship run, the Wildcats won 23 games and had 14 ties.